### v1

# Risk-inclusive Contextual Bandits for Early Phase Clinical Trials

Abstract
Early-phase clinical trials face the challenge of selecting optimal drug doses that balance safety and efficacy due to uncertain dose-response relationships and varied participant characteristics. Traditional randomized dose allocation often exposes participants to sub-optimal doses by not considering individual covariates, necessitating larger sample sizes and prolonging drug development. This paper introduces a risk-inclusive contextual bandit algorithm that utilizes multi-arm bandit (MAB) strategies to optimize dosing through participant-specific data integration. By combining two separate Thompson samplers, one for efficacy and one for safety, the algorithm enhances the balance between efficacy and safety in dose allocation. The effect sizes are estimated with a generalized version of asymptotic confidence sequences (AsympCS) (Waudby-Smith et al.,, 2024), offering a uniform coverage guarantee for sequential causal inference over time. The validity of AsympCS is also established in the MAB setup with a possibly mis-specified model. The empirical results demonstrate the strengths of this method in optimizing dose allocation compared to randomized allocations and traditional contextual bandits focused solely on efficacy. Moreover, an application on real data generated from a recent Phase IIb study aligns with actual findings.
Keywords: Anytime-valid inference; Dose-ranging studies; Efficacy and safety; Model-assisted inference; Response adaptive randomization.
1 Introduction
Randomized Controlled Trials (RCTs) are structured studies to evaluate the safety and effectiveness of investigational treatments. RCTs are conducted in progressive phases (Friedman et al.,, 2010). Phase I trials involve a small group to assess safety and tolerability. Phase II expands the sample size to evaluate preliminary efficacy and safety, establishing optimal dosage. This precedes larger Phase III trials, which confirm treatment safety and effectiveness. Of particular importance are dose-ranging studies, which are frequent in Phase IIb and aim to determine the optimal dose that balances efficacy and safety for further development. These trials administer multiple doses to participant cohorts to identify the best therapeutic benefit with minimal adverse effects, as illustrated in Figure 1.
Figure 1: Study design of a Phase II/III clinical trial. In this scenario, Dose 5 of a novel drug, demonstrating preliminary safety and early efficacy in Phase I and IIa, progresses to Phase IIb. Then, five doses and a placebo control are randomized in Phase IIb to facilitate dose-response curve characterization and optimal dose selection. Final analyses are performed at the end of Phase IIb to select the most suitable dose for the larger Phase III trial.
Traditional dose-ranging study designs often rely on equal randomization (ER). However, ER has several limitations (Hay et al.,, 2014): (1) The unknown dose-response relationship in early phases exposes many participants to suboptimal doses. (2) Dose-response relationships are often heterogeneous and influenced by participant-level covariates (Le Tourneau et al.,, 2009). Adaptive methods such as response-adaptive randomization (RAR) (Robertson et al.,, 2023) and covariate-adjusted response-adaptive (CARA) designs (Villar and Rosenberger,, 2018; Yang et al.,, 2024) aim to address this; nevertheless, existing adaptive designs fall short in combining efficacy, safety, and covariates to inform treatment allocation. (3) Traditional approaches only allow formal inference at the end, whereas continuous monitoring is needed to enable early stopping for efficacy or futility, saving resources and time. (4) Finally, a valid inference on effect sizes often depends on correctly specified parametric models, but practitioners often do not have substantial knowledge about the efficacy or safety in early phases, incurring model mis-specification. As early-phase trials involve small sample sizes and investigational drugs that are not yet approved, practitioners prefer inference methods that do not rely heavily on model assumptions, especially given the critical role such results play in informing Phase III planning and decision-making. Overcoming these limitations is crucial for accelerating drug development while ensuring participant safety.
Motivated by these challenges, we propose an RAR framework extending recent developments in online reinforcement learning (RL, Agrawal and Goyal,, 2013) and sequential causal inference (Waudby-Smith et al.,, 2024) to optimize dose-ranging trials. Our framework has two components: trial optimization and trial evaluation. For optimization, we frame the trial as a contextual multi-arm bandit (CMAB) problem, incorporating covariate information. To balance efficacy and safety, we combine two separate CMAB models, one for each, with a pre-specified weight parameter. The weight parameter enables informed decision-making based on both trial data and insights from preclinical and earlier-phase studies. Trial evaluation focuses on estimating effect sizes for all active doses. Moving beyond standard post-trial inference, we perform anytime-valid inference by extending Asymptotic Confidence Sequences (AsympCS, Waudby-Smith et al.,, 2024) to the CMAB setup, offering coverage guarantees during continuous data monitoring and allowing statistically principled early trial stopping. While parametric models are used for efficient trial optimization, inference during trial evaluation is robust to model mis-specification, making our approach model-assisted.
Extensive simulations and analysis based on a real-life Phase IIb study demonstrate the utility and effectiveness of the proposed approach. Notably, in our experiments, we also consider some logistical challenges to operate the proposed method in a multi-center, multi-regional setup. We further discuss unaddressed issues in Section 5.
The rest of this article is organized as follows: Section 1.1 reviews relevant work; Section 2 defines notations, introduces the setup, and details our proposed method; Section 3 examines its operating characteristics on simulated data; Section 4 presents its application to a real-world dataset; and Section 5 concludes by discussing future directions.
1.1 Related Work
Adaptive Designs in Clinical Trials
Adaptive clinical trials offer advantages over traditional fixed designs by enabling modifications based on accumulating data (Pallmann et al.,, 2018; FDA,, 2019; Burnett et al.,, 2020). However, adaptive designs, including Response Adaptive Randomization (RAR), may perform sub-optimally in long-duration trials where treatment effects evolve over time (Proschan and Evans,, 2020). Classical RAR methods are often criticized for reacting too quickly to certain trends in the data and for ignoring the covariates of participants during arm allocation. Most existing methods fall short of analyzing data from adaptive experiments (Thall et al.,, 2015). Despite these challenges, the advantages of adaptive designs have motivated methodological research in the past decade (e.g., Villar et al.,, 2015; Williamson and Villar,, 2020; Robertson et al.,, 2023; Norwood et al.,, 2024). In this article, we focus on short-duration trials with approximately constant treatment effect sizes, common in areas such as oncology and dermatology. Our method addresses the foregoing issues in classical adaptive designs by (1) making appropriate modifications, such as “clipping,” to encourage exploring sub-optimal arms, (2) using participants’ covariate information during arm allocation, and (3) performing sequential causal inference that allows valid model-robust inference on effect sizes.
Safety Consideration and Multi-objective Reinforcement Learning
RAR methods that balance efficacy and safety remain underexplored. To address this gap, we utilize Thompson Sampling (TS, Thompson,, 1933) and leverage recent developments in online RL (Agrawal and Goyal,, 2013). Although multi-objective contextual bandit methods allow integration of multiple responses (e.g., Kim et al.,, 2024), existing methods typically treat all outcomes equally. Clinical trials, however, often have unique safety tolerance limits for efficacy, such as in oncology, where greater risks might be accepted for higher therapeutic benefits, versus dermatology, where minimal adverse effects are preferred. Furthermore, by Phase IIb, preliminary safety knowledge usually exists, warranting its incorporation. Although weighted composite outcomes are used in precision medicine (Luckett et al.,, 2021; Zhu et al.,, 2024), existing work focuses on individualized treatment recommendations. To our knowledge, no existing work effectively combines both efficacy and safety endpoints to develop multi-objective contextual bandits for early-phase clinical trials.
Statistical Inference from Reinforcement Learning
Research in online RL mainly focuses on regret minimization, balancing exploration and exploitation (Agrawal and Goyal,, 2017; Dimakopoulou et al.,, 2019, 2021; Kim et al.,, 2021). However, all phases of clinical trials typically involve smaller sample sizes than online RL literature often considers. In small samples, some suboptimal arm allocations are still manageable due to exploration, with valid estimation and inference of estimands like effect sizes being the primary interest. Recent inference tools enable us to perform valid inference with adaptively collected data (Bibaut et al.,, 2021; Hadad et al.,, 2021; Zhang et al.,, 2020, 2021; Shen et al.,, 2023). While these methods primarily focus on post-trial inference, continuous monitoring in early-phase trials facilitates faster progression to Phase III or early termination when the investigational drug is evidently effective or deemed unsafe, helping reduce unnecessary time and resource expenditure. Valid sequential testing offers this flexibility, finding success in other domains (Lindon et al.,, 2022; Maharaj et al.,, 2023). More recently, Waudby-Smith et al., (2024) proposed AsympCS for time-uniform inference of effect sizes in both randomized and observational data. Based on this, Dalal et al., (2024) extended its validity for general causal parameters. Additionally, Ham et al., (2022) proposed a design-based confidence sequence for anytime-valid inference of finite-sample estimands. While non-asymptotic CSs (Howard et al.,, 2021; Waudby-Smith and Ramdas,, 2023) are popular, we advocate for AsympCS because it yields narrower confidence bounds in small sample sizes.
2 Methods
2.1 Framework
Our setup for dose-ranging trials includes $K$ arms. Arm 111 corresponds to placebo, and Arms $2,\ldots,K$ are active doses of a novel drug with increasing order of doses. For each participant, we consider their covariate information and two (continuous) endpoints, efficacy and safety. Specifically, assume the complete profiles of participants $(\bm{X}_{n},\{R_{n}(a)\}_{a=1}^{K},\{S_{n}(a)\}_{a=1}^{K})_{n\geq 1}$ are independent and identically distributed, where $\bm{X}_{n}\in\mathbb{R}^{d}$, $\{R_{n}(a)\}_{a=1}^{K}$, and $\{S_{n}(a)\}_{a=1}^{K}$ are the $n$-th participant’s covariate vector, potential outcomes of efficacy and safety, respectively. The complete profile of each participant cannot be observed. Thus, we randomly assign Arm $A_{n}$ to the $n$-th participant, and observe the outcomes $R_{n}$ and $S_{n}$ instead. For equal randomization (ER), $A_{n}$ is independent of the covariates $\bm{X}_{n}$ and other participants’ information. Response adaptive randomization (RAR) improves arm allocation by allowing the distribution of $A_{n}$ to depend on the covariates $\bm{X}_{n}$ and the historical information $\mathcal{H}_{n-1}=\{(\bm{X}_{i},A_{i},R_{i},S_{i})\}_{i=1}^{n-1}$ from the first $n-1$ participants. In this article, we focus on identifying the most efficacious arm with uncertainty quantification, while allocating better and safer treatments to participants. Define the average efficacy of Arm $a$ as $\theta(a):=\mathbb{E}[R_{n}(a)]$ and the effect size of Arm $a$ relative to the placebo as $\Delta(a):=\theta(a)-\theta(1)$. Average safety can also be defined in a similar manner. For ease of exposition, we focus on average efficacy in this article. The following conditions suffice to identify $\{\theta(a)\}_{a=1}^{K}$ and $\{\Delta(a)\}_{a=2}^{K}$:
  1. (A1)
    Consistency: $R_{n}=\sum_{a=1}^{K}R_{n}(a)\mathbb{I}\{A_{n}=a\}$ and $S_{n}=\sum_{a=1}^{K}S_{n}(a)\mathbb{I}\{A_{n}=a\}$; $n\geq 1$.
  2. (A2)
    Positivity: $\mathbb{P}\left[{A_{n}=a|\bm{X}_{n},\mathcal{H}_{n-1}}\right]>0$ for all $a\in[K]$; $n\geq 1$.
  3. (A3)
    Ignorability: $\{R_{n}(a),S_{n}(a)\}_{a=1}^{K}\perp\!\!\!\!\perp A_{n}|\{\bm{X}_{n},\mathcal{H}_{n-1}\}$; $n\geq 1$.
By (A1), we assume participants adhere to their assigned treatments. We will design a trial optimization algorithm in Section 2.2, so that (A2)–(A3) are automatically met. It further leads to trial evaluation through inferring $\{\Delta(a)\}_{a=2}^{K}$ in Section 2.3.
To enable a Bayesian trial optimization, we introduce a parametric working model for the potential efficacy and safety outcomes:
$R_{n}(a)=\mu^{(a)}(\bm{X}_{n})+\epsilon_{n},\quad S_{n}(a)=\nu^{(a)}(\bm{X}_{n})+\delta_{n},\quad n\geq 1,$ (1)
where $\mu^{(a)}(\bm{x})=\beta_{a,0}+\bm{x}^{T}\bm{\beta}_{a}$, $\nu^{(a)}(\bm{x})=\gamma_{a,0}+\bm{x}^{T}\bm{\gamma}_{a}$, $\epsilon_{n}\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$ and $\delta_{n}\sim\mathcal{N}(0,\sigma_{\delta}^{2})$ are normal errors. We will show that parametric model (1) assists in efficient learning for better treatment allocation. More importantly, the validity of inference for $\{\Delta(a)\}_{a=2}^{K}$ will not be affected even if (1) is mis-specified.
2.2 Trial Optimization using a Risk-inclusive Thompson Sampler
An Overview of TS
Under consistency (A1) and parametric form (1), $R_{i}=\beta_{A_{i},0}+\bm{X}_{i}^{T}\bm{\beta}_{A_{i}}^{*}+\epsilon_{i}.$ TS starts with priors $\{\pi^{(0)}(\bm{\beta}_{a})\}_{a=1}^{K}$ on the regression parameters, updating them to posteriors $\{\pi^{(n)}(\bm{\beta}_{a})\}_{a=1}^{K}$ based on historical data from the first $n-1$ participants. Before assigning arm to the $n$-th participant, TS simulates $MK$ samples, with $M$ samples drawn from each posterior in $\{\pi^{(n)}(\bm{\beta}_{a})\}_{a=1}^{K}$. Let $\{\bm{b}^{(n)}_{m}(a)\}_{m=1}^{M}$ be the $M$ samples from $\pi^{(n)}(\bm{\beta}_{a})$. The next arm is assigned according to $\{(p_{n}(a,\bm{x})\}_{a=1}^{K}$, where $p_{n}(a,\bm{x}):=\frac{1}{M}\sum_{m=1}^{M}\mathbb{I}\{\bm{x}^{T}\bm{b}^{(n)}_{m}(a)\in\underset{a^{\prime}\in[K]}{\operatorname*{arg\,max}}\ \bm{x}^{T}\bm{b}^{(n)}_{m}(a^{\prime})\}$ is the propensity of allocating arm $a$ to the $n$-th participant, given their covariate $\bm{X}_{n}=\bm{x}$ and the historical data $\mathcal{H}_{n-1}=\{(\bm{X}_{i},A_{i},R_{i},S_{i})\}_{i=1}^{n-1}$. Safety outcomes $\{S_{i}\}_{i=1}^{n-1}$, although available in $\mathcal{H}_{n-1}$, are ignored in this approach.
Outline of RiTS
To utilize and balance safety, we propose using two separate TS: one for efficacy and one for safety. The priors $\{\pi^{(0)}(\bm{\beta}_{a})\}_{a=1}^{K}$ for efficacy, and $\{\pi^{(0)}(\bm{\gamma}_{a})\}_{a=1}^{K}$ for safety are updated to posteriors $\{\pi^{(n)}(\bm{\beta}_{a})\}_{a=1}^{K}$ and $\{\pi^{(n)}(\bm{\gamma}_{a})\}_{a=1}^{K}$ using $\mathcal{H}_{n-1}$. For each arm, samples $\{\bm{b}_{m}^{(n)}(a)\}_{m=1}^{M}$ and $\{\bm{g}_{m}^{(n)}(a)\}_{m=1}^{M}$ are simulated from $\pi^{(n)}(\bm{\beta}_{a})$ and $\pi^{(n)}(\bm{\gamma}_{a})$, respectively. For a weight parameter $w\in(0,1)$, define $\omega(\bm{x},\bm{b},\bm{g};w):=w\cdot\bm{x}^{T}\bm{b}+(1-w)\cdot\bm{x}^{T}\bm{g}$. The next arm is randomized according to $\{q_{n}(a,\bm{x})\}_{a=1}^{K}$, where given $\bm{X}_{n}=\bm{x}$ and past data $\mathcal{H}_{n-1}$,
$q_{n}(a,\bm{x}):=\frac{1}{M}\sum_{m=1}^{M}\mathbb{I}\left\{\omega(\bm{x},\bm{b}^{(n)}_{m}(a),\bm{g}^{(n)}_{m}(a);w)\in\operatorname*{arg\,max}_{a^{\prime}\in[K]}\omega(\bm{x},\bm{b}^{(n)}_{m}(a^{\prime}),\bm{g}^{(n)}_{m}(a^{\prime});w)\right\}$
is the propensity of assigning arm $a$ to the $n$-th participant. We refer to this approach as the “Risk-inclusive Thompson Sampling” (RiTS).
Practical Considerations to Choose Weight Parameter $w$
[…]
Prior Distributions
Multivariate normal priors are chosen due to the availability of closed-form posteriors in our setup. More informative priors can be incorporated based on domain knowledge. For intractable posteriors, a variational approximation can be performed. Bayesian regression serves as an online regression algorithm, where domain knowledge from previous trials can be easily incorporated.
Exact Form of Posteriors
Let $\mathcal{N}_{d}(\bm{\mu},\mathbf{\Omega}^{-1})$ be the $d$-variate normal distribution with mean $\bm{\mu}$ and precision matrix $\mathbf{\Omega}$. For all $n\geq 1$, let $\pi^{(n)}(\bm{\beta}_{a}):=\mathcal{N}_{d}\big{(}\bm{\mu}_{a}^{(n)},\{\mathbf{\Omega}_{a}^{(n)}\}^{-1}\big{)}$. Fix $\bm{\mu}_{a}^{(0)}:=\bm{0}$, $\mathbf{\Omega}_{a}^{(0)}:=\mathbf{I}$. By the conjugacy of normal priors, update parameters by $\mathbf{\Omega}_{a}^{(n)}:=\mathbf{\Omega}_{a}^{(n-1)}+\mathbb{I}\{A_{n-1}=a\}\frac{1}{\sigma_{0}^{2}}\bm{X}_{n-1}\bm{X}_{n-1}^{T}$ and $\bm{\mu}_{a}^{(n)}:=\{\mathbf{\Omega}_{a}^{(n)}\}^{-1}\big{[}\mathbb{I}\{A_{n-1}=a\}\frac{1}{\sigma_{0}^{2}}\bm{X}_{n-1}R_{n-1}+\mathbf{\Omega}_{a}^{(n-1)}\bm{\mu}_{a}^{(n-1)}\big{]}$ for all $n\geq 1,\ a\in[K]$. Similar steps update the posteriors for safety, i.e., $\{\pi^{(n)}(\bm{\gamma}_{a})\}_{a=1}^{K}$.
2.3 Robust Trial Evaluation Using AsympCS
Construction of AsympCS
For $n\geq 1$, let $Z_{n}:=(\bm{X}_{n},A_{n},R_{n},S_{n})$. To begin, we randomly allocate arms to the first $n_{0}$ participants using equal randomization. Once we observe $\{Z_{i}\}_{i=1}^{n_{0}}$, arms are randomized adaptively using RiTS to the $(n_{0}+1)$-th participant onward. Odd and even-indexed $Z_{i}$ are placed in $\mathcal{D}^{trn}_{n}$ and $\mathcal{D}^{eval}_{n}$ respectively, i.e., $\mathcal{D}^{trn}_{n}:=(Z_{1},Z_{3},\ldots,Z_{\left\lceil{n/2}\right\rceil})$ and $\mathcal{D}^{eval}_{n}=(Z_{2},Z_{4},\ldots,Z_{\left\lfloor{n/2}\right\rfloor})$. Denote $N^{\prime}:=|\mathcal{D}^{trn}_{n}|=\left\lceil{n/2}\right\rceil$ and $N:=|\mathcal{D}^{eval}_{n}|=\left\lfloor{n/2}\right\rfloor$. Corresponding sets of indicators are defined as $\mathcal{I}^{trn}_{n}$ and $\mathcal{I}^{eval}_{n}$, i.e., $i\in\mathcal{I}^{trn}_{n}$ if and only if $Z_{i}\in\mathcal{D}^{trn}_{n}$. With $\{Z_{i}\}_{i=1}^{n}$, we construct $(1-\alpha)$ AsympCS $\{\mathcal{C}_{n}^{(\alpha)}(a)\}_{n\geq 1}$ for $\Delta(a)$ using the four steps below:
  1. (Step 1)
    Regression: Obtain regression estimator $\hat{\mu}_{N^{\prime}}^{(a)}(\bm{x})$ for $\mathbb{E}(R|A=a,\bm{X}=\bm{x})$ using $\mathcal{D}^{trn}_{n}$ for all $a\in[K]$.
  2. (Step 2)
    Pseudo-outcomes: For each $i\in\mathcal{I}^{eval}_{n}$, compute the Augmented Inverse Propensity Weighted (AIPW) pseudo-outcomes for efficacy endpoint:
    $\hat{g}_{N^{\prime}}(Z_{i};a):=\hat{\mu}_{N^{\prime}}^{(a)}(\bm{X}_{i})+\frac{\mathbb{I}\{A_{i}=a\}}{q_{i}(a,\bm{X}_{i})}\left\{{R_{i}-\hat{\mu}_{N^{\prime}}^{(A_{i})}(\bm{X}_{i})}\right\},\quad\forall a\in[K],$
    and pseudo-contrasts $\hat{f}_{N^{\prime}}(Z_{i};a)=\hat{g}_{N^{\prime}}(Z_{i};a)-\hat{g}_{N^{\prime}}(Z_{i};1)$ for $a\in[K]/\{1\}$.
  3. (Step 3)
    Cross-fitting: Steps 1–2 are replicated by switching the roles of $\mathcal{D}^{trn}_{n}$ and $\mathcal{D}^{eval}_{n}$ to obtain a tuple of pseudo-contrasts for all observed efficacy endpoints, i.e., for each $R_{i}$, generate $\{\hat{f}_{N^{\prime}}(Z_{i};a)\}_{a=2}^{K}$ or $\{\hat{f}_{N}(Z_{i};a)\}_{a=2}^{K}$ depending on $i\in\mathcal{I}^{eval}_{n}$ or $\mathcal{I}^{trn}_{n}$.
  4. (Step 4)
    Construction of $\mathcal{C}_{n}^{(\alpha)}(a)$: Let $\hat{\Delta}_{n}^{\times}(a):=\big{(}\sum_{i\in\mathcal{I}^{eval}_{n}}\hat{f}_{N^{\prime}}(Z_{i};a)+\sum_{i\in\mathcal{I}^{trn}_{n}}\hat{f}_{N}(Z_{i};a)\big{)}/n$, and $\hat{\sigma}_{n}^{2}(a):=\big{(}\hat{{\rm var}}_{N}(\hat{f}_{N^{\prime}};a)+\hat{{\rm var}}_{N^{\prime}}(\hat{f}_{N};a)\big{)}/2$, where $\hat{{\rm var}}_{N}(\hat{f}_{N^{\prime}};a)$ and $\hat{{\rm var}}_{N^{\prime}}(\hat{f}_{N};a)$ are the sample variances of $\{\hat{f}_{N^{\prime}}(Z_{i};a):i\in\mathcal{I}^{eval}_{n}\}$ and $\{\hat{f}_{N}(Z_{i};a):i\in\mathcal{I}^{trn}_{n}\}$, respectively. Obtain the $n$-th interval of the AsympCS:
    $\mathcal{C}_{n}^{(\alpha)}(a):=\hat{\Delta}_{n}^{\times}(a)\pm\sqrt{{2(n\rho^{2}\hat{\sigma}_{n}^{2}(a)+1)\log\left({\frac{\sqrt{n\rho^{2}\hat{\sigma}_{n}^{2}(a)+1}}{\alpha}}\right)}/{n^{2}\rho^{2}}}.$ (2)
    Here, $\rho>0$ is a tuning parameter and the asymptotic time-uniform guarantee of $\{\mathcal{C}_{n}^{(\alpha)}(a)\}_{n\geq m}$ holds for any $\rho>0$, provided the burn-in sample size $m$ is sufficiently large. We follow the recommendation of Waudby-Smith et al., (2024) and fix $\rho=\rho_{m}:=\sqrt{\frac{-2\log\alpha+\log(-2\log\alpha)+1}{\hat{\sigma}_{m}(a)m\log(\max\{m,e\})}}$ where $e$ is the natural base. The choice of $\rho=\rho_{m}$ ensures $\mathcal{C}_{m}^{(\alpha)}(a)$ is the narrowest.
Our proposed method is summarized in Algorithm 1. For implementation, we use a weighted ridge regression estimator $\hat{\mu}_{N^{\prime}}^{(a)}(\bm{x}):=\hat{\beta}_{a,0}+\bm{x}^{T}\hat{\bm{\beta}}$ in Step 1, where for tuning parameter $\lambda>0$,
$\{\hat{\beta}_{a,0}\},\ \hat{\bm{\beta}}:=\operatorname*{arg\,min}_{\{\beta_{a,0}\},\ \bm{\beta}}\left[{\sum_{a\in[K]}\sum_{i\in\mathcal{I}^{trn}_{n}}\frac{\mathbb{I}\{A_{i}=a\}}{q_{i}(a,\bm{X}_{i})}(R_{i}-\beta_{a,0}-\bm{X}_{i}^{T}\bm{\beta})^{2}+\lambda\left\|{\bm{\beta}}\right\|_{2}^{2}}\right].$ (3)
A more flexible model can be used, but requires a larger burn-in sample size. It is noteworthy that the inverse propensity weight adjustment is crucial for valid coverage of AsympCS when the model is possibly mis-specified and the data are adaptively collected; see Section 2.4.
Algorithm 1 Proposed Framework of RiTS
1:Data $\{\bm{X}_{n}\}_{n\geq 1}$; Priors $\pi(\bm{\beta}_{a}),\ \pi(\bm{\gamma}_{a})$; Weight $w\in[0,1]$; Number of initial participants to be randomized $n_{0}$; Minimum propensity for arms (clipping parameter) $\delta\in(0,1/K)$; Number of burn-in samples $m$.
2:for $n\leq n_{0}$ do
3:  Randomize arms; Set $q_{n}(a,\bm{X}_{n})=1/K$ for all $a\in[K]$.
4:  Simulate $A_{n}\sim\text{Multinomial}\left[{\left\{{q_{n}(a,\bm{X}_{n})}\right\}_{a=1}^{K}}\right]$ and assign arm $A_{n}$ to $n$-th participant.
5:end for
6:for $n\geq n_{0}+1$ do
7:  Utilize $\mathcal{H}_{n-1}$ to update priors to posteriors $\{\pi^{(n)}(\bm{\beta}_{a})\}_{a=1}^{K}$, $\{\pi^{(n)}(\bm{\gamma}_{a})\}_{a=1}^{K}$.
8:  Obtain $\{q_{n}(a,\bm{X}_{n})\}_{a=1}^{K}$; Clip to ensure $q_{n}(a,\bm{X}_{n})\in[\delta,1-\delta]$ for $a\in[K]$.
9:  Simulate $A_{n}\sim\text{Multinomial}\left[{\left\{{q_{n}(a,\bm{X}_{n})}\right\}_{a=1}^{K}}\right]$ and assign arm $A_{n}$ to $n$-th participant.
10:  if $n\geq m$ then
11:   Use (2) to obtain $\{\mathcal{C}_{n}^{(\alpha)}(a)\}_{a=2}^{K}$.
12:  end if
13:end forreturn $\{(\bm{X}_{n},A_{n},R_{n},S_{n})\}_{n\geq 1}$ and $\{\mathcal{C}_{n}^{(\alpha)}(a):a\in\{2,\ldots,K\}\}_{n\geq m}$.
Hyper-parameters in AsympCS
All hyper-parameters in Algorithm 1 need to be chosen carefully so that AsympCS controls the cumulative miscoverage at a nominal level. The asymptotic validity of AsympCS is tied to $n_{0}$, which has to increase with the increase in the expected total sample size of the trial. Fixing $n_{0}$ to a smaller number and $\delta$ too close to zero leads to an inadequate number of participants in potentially sub-optimal arms. Without any regularization, individual regression estimators may suffer from over-fitting. In all of our experiments, we use the ridge penalty to avoid such issues. Even with regularization, AsympCS fails to control the cumulative miscoverage after fixing both $n_{0}$ and $\delta$ to smaller values. Web Appendix A.2 explores different choices of hyperparameters and their impact on the empirical cumulative miscoverage.
Batched Training and Delayed Response for Practical Implementation
Delayed response and batched training are also allowed in Algorithm 1. Specifically, Line 6 and Line 10 of Algorithm 1 can be executed with potentially delayed responses $(r_{i},s_{i})$ that are available in $\mathcal{D}^{trn}_{n}$ and $\mathcal{D}^{eval}_{n}$. Batched training is also possible by implementing Line 6 of Algorithm 1 for batches of observations, as incorporated in our R code.
2.4 Validity of AsympCS
The validity of AsympCS hinges upon Assumptions 12. For any function $f(\cdot)$ from $\mathbb{R}^{d}$ to $\mathbb{R}$, define $\|f\|_{L_{2}}:=(\mathbb{E}_{\bm{X}}|f(\bm{X})|^{2})^{1/2}$, where $\mathbb{E}_{X}(\cdot)$ is taken with respect to $\bm{X}$.
Assumption 1
For $a\in[K]$, $\|\hat{\mu}_{n}^{(a)}-\bar{\mu}^{(a)}\|_{L_{2}}=o(1)$ a.s., for some function $\bar{\mu}^{(a)}$.
Assumption 2
There exists $\eta>0$ such that $\max_{a\in[K]}\|\hat{\mu}_{n}^{(a)}\|_{L_{2+\eta}}\vee\|\bar{\mu}^{(a)}\|_{L_{2+\eta}}=O(1)$ a.s.
Theorem 1
For Algorithm 1, the following statements are true:
  • Conditions (A2)–(A3) in Section 2.1 hold by design.
  • Under (A1), if $\{\bm{X}_{n}\}_{n\geq 1}$ are sub-Gaussian and the estimator (3) are used to construct $\{\mathcal{C}_{n}^{(\alpha)}(a)\}_{n\geq 1}$, then Assumptions 12 hold.
  • Under (A1), if Assumptions 12 hold, then $\{\mathcal{C}_{n}^{(\alpha)}(a)\}_{n\geq 1}$ is a $(1-\alpha)$-AsympCS for $\Delta(a)$.
Theorem 1 only requires the regression estimators $\{\hat{\mu}^{(a)}_{n}\}_{a=1}^{K}$ converge uniformly to fixed functions $\{\bar{\mu}^{(a)}\}_{a=1}^{K}$, which is not necessarily the true functions $\{\mu^{(a)}\}_{a=1}^{K}$. In practice, it is often difficult to model the efficacy and safety correctly. This flexibility makes our framework model-assisted: the parametric model (1) assists in trial optimization and estimation of effect sizes, but mis-specifying (1) does not void inferential results from AsympCS. Nevertheless, the existence of $\{\bar{\mu}^{(a)}\}_{a=1}^{K}$ relies on the choice of estimator, especially when the model is mis-specified and the data are adaptively collected. The inverse propensity weighting in (3) ensures the existence of $\{\bar{\mu}^{(a)}\}_{a=1}^{K}$, while without this adjustment, $\{\bar{\mu}^{(a)}\}_{a=1}^{K}$ may not exist at all. Moreover, enough participants have to be allocated to each arm; otherwise, $\hat{\mu}^{(a)}_{n}$ may not be sufficiently convergent to $\bar{\mu}^{(a)}$ uniformly. Fixing $n_{0},\delta$, and $m$ to moderate values improves this uniform convergence with a relatively small sample size.
We also note that Theorem 6 in Cook et al., (2024) is not applicable here, as their proof does not incorporate estimation of the efficient influence function (EIF) using observed data.
3 Simulation Studies
The experiment is structured using the ADEMP setup (Morris et al.,, 2019). A total of $n_{sim}=1000$ replications are performed on simulated datasets of size $n_{obs}=200$.
Aims
Our goal is two-fold: (1) Find “the best arm” with confidence as quickly as possible, (2) Allocate more participants to better treatments by balancing efficacy and safety. Note that, while allocation ratios indicate relative arm ordering, it hinges upon correct specifications of $\{\mu^{(a)}(\bm{x})\}_{a=1}^{K}$ and $\{\nu^{(a)}(\bm{x})\}_{a=1}^{K}$. Therefore, the best arm should not be identified solely by the allocation ratio. Instead, it is typically identified based on both estimated efficacy and safety, which should be robust to model mis-specifications. For simplicity, we focus only on the efficacy endpoint and use “the best arm” and “the most efficacious arm” interchangeably. Depending on $w$, AsympCS for average utility can also be used instead to identify the best arm for Phase III.
Data-generating Mechanisms
Early-phase clinical trials typically involve a limited cohort of participants. Although many demographic covariates and assays are measured, drug developers generally focus on a small subset of participant-level covariates to assess heterogeneous dose responses. Key covariates are often chosen based on domain knowledge from previous studies. Following standard practice, we consider a single participant-level covariate, denoted by $Z$, which is simulated from a standard normal distribution, $Z\sim\mathcal{N}(0,1)$.
We first consider a high signal-to-noise (High-SNR) scenario, where the true average efficacy as a function of $Z$ across each arm is fixed as follows, $\widetilde{\mu}_{0}(z,1)=2-0.01(z+1/2)^{2}-0.01(z-1/2)^{2}$, $\widetilde{\mu}_{0}(z,2)=2.7-0.2(z-1/2)^{2}-0.2(z+1/2)^{2}$, $\widetilde{\mu}_{0}(z,3)=2.7-0.01(z-1/2)^{2}-0.01(z+1/2)^{2}$ and $\widetilde{\mu}_{0}(z,4)=3.2-0.2(z-1/2)^{2}-0.2(z+1/2)^{2}$. True effect sizes are $\bm{\Delta}:=(\Delta(2),\Delta(3),\Delta(4))=(0.225,0.475,0.725)$. The overall effects are calculated by taking expectations of $\widetilde{\mu}_{0}(Z,a)$ with respect to $Z\sim\mathcal{N}(0,1)$. The true average safety across each arm is also a quadratic function of one-dimensional context with $\widetilde{\nu}_{0}(z,1)=2$, $\widetilde{\nu}_{0}(z,2)=2-0.01z^{2}$, $\widetilde{\nu}_{0}(z,3)=2-0.1z^{2}$ and $\widetilde{\nu}_{0}(z,4)=2-0.6z^{2}$. Define $\mu_{0}(\bm{x},a):=\widetilde{\mu}_{0}(z,a)$, where $\bm{x}=(z,z^{2})^{T}$ is the covariate for a participant. Therefore, the true generating distributions for efficacy and safety are as follows: $R_{i}=\mu_{0}(\bm{X}_{i},A_{i})+\epsilon_{i}$, $S_{i}=\nu_{0}(\bm{X}_{i},A_{i})+\delta_{i}$, where $\epsilon_{i},\delta_{i}\sim\mathcal{N}(0,1)$ are independent. Next, a low signal-to-noise (Low-SNR) setup is also considered where $\mu_{0}(\cdot,\cdot)$ and $\nu_{0}(\cdot,\cdot)$ are replaced by $\mu_{0}(\cdot,\cdot)/2$ and $\nu_{0}(\cdot,\cdot)/2$, respectively. Effect sizes are $\bm{\Delta}=(0.1125,0.2375,0.3625)$ in the Low-SNR setup.
In the above setup, Arm 1 represents the placebo, while Arms 2–4 correspond to the drug’s low, medium, and high doses, respectively. For simplicity, we define the average utility as a weighted combination of efficacy and safety with equal weights: $\upsilon_{0}(z,a)=(1/2)\mu_{0}(z,a)+(1/2)\nu_{0}(z,a)$. Arm 1 is the safest but has the lowest efficacy, resulting in the lowest utility. Arm 2 consistently exhibits lower efficacy, safety, and utility across all participants, and adaptive trials are expected to allocate very few participants to this arm. Dose-ranging trials often include a low-dose arm to better characterize the dose-response curve. Arm 2 serves this purpose here. Arm 4 provides greater efficacy than Arm 3 for a small group of participants with $z$ values near 0, while Arm 3 is the overall optimal arm when considering both efficacy and safety. However, efficacy is the primary criterion for selecting the best dose for Phase III, and therefore, we expect to choose Arm 4 as the best arm. Alternative criteria exist for selecting the best dose for Phase III, which consider both efficacy and safety. For simplicity, we consider the best dose as the most efficacious arm in our simulation experiments.
Estimand/target of Analysis
The primary target is to estimate $\{\Delta(a)\}_{a=2}^{K}$ and identify (with confidence) the arm having the largest $\Delta(a)$ as quickly as possible. We use AsympCSs with a stopping criterion: A trial is stopped once one of the confidence sequences is above 0.10.10.1, the minimum clinically significant effect size, and the best arm is chosen based on the highest estimated effect size. The secondary target is to minimize cumulative regrets (to be defined later), which monitors the allocation of better arms in efficacy, safety, and utility.
Methods
We compare the performance of our proposed method, referred to as RiTS, with one Thompson Sampler trained solely on the efficacy endpoint, referred to as TS. We also include a randomized allocation method, referred to as Rand, as a baseline. For each replication, we first simulate a dataset of size $n_{sim}$. Then, the first $n_{0}$ participants receive a randomly selected arm. Since endpoints are expected to be observed with some delay, we consider a delay parameter and fix it to 101010. We utilize observations from the first $n_{0}-\texttt{delay}$ participant and utilize one of Rand, TS and RiTS to allocate an arm to the $(n_{0}+1)$-th participant onward. Such delay is continued for all future allocations, i.e., to allocate an arm to the $n$-th participant, we only utilize observations from the first $(n-\texttt{delay})$ participants. The trial evaluation compares the performance of AsympCS with a current baseline of two-sample t-tests (T-test, hereafter) and an inverse propensity weighted (IPW) estimator in AsympCS. The last method is equivalent to setting $\{\hat{\mu}_{n}^{(a)}(\cdot)\}_{a=1}^{K}$ to zero for all arms. To assess the asymptotic behavior, we start the construction of AsympCS with a full observation from the first $m$ participant. T-test is only included for Rand allocation (Rand-T-test, hereafter) as T-test is well-known to be inadequate for adaptively collected data. In what follows, Rand-AIPW, TS-AIPW, and RiTS-AIPW refer to the original AsympCS method with the corresponding method. Rand-IPW, TS-IPW and RiTS-IPW refer to the modified version of AsympCS after replacing the AIPW estimator with the IPW estimator.
Performance Measures
To assess the quality of AsympCS, we monitor bias, root mean squared error (RMSE), width, and cumulative miscoverage rate. We also include metrics such as the proportion of replications where Arm 4 has the highest estimated effect size (‘Winner (Arm 4)’ or WA4, hereafter), and where the stopping criterion is met (‘Stopping Criteria’ or SC, hereafter). Note that WA4 only uses point estimates, whereas SC only uses bounds of AsympCS. For our secondary target, cumulative regrets are monitored, defined as follows. For the $i$-th participant with covariate $\bm{x}_{i}$, define the optimal arm based on utility, efficacy, and safety, respectively, as $a_{i}^{*}:=\operatorname*{arg\,max}_{a\in\mathcal{A}}\bm{x}_{i}^{T}[w\bm{\beta}_{a}^{*}+(1-w)\bm{\gamma}_{a}^{*}]$, $a_{i}^{(e)*}:=\operatorname*{arg\,max}_{a\in\mathcal{A}}\bm{x}_{i}^{T}\bm{\beta}_{a}^{*}$ and $a_{i}^{(s)*}:=\operatorname*{arg\,max}_{a\in\mathcal{A}}\bm{x}_{i}^{T}\bm{\gamma}_{a}^{*}$. Corresponding cumulative regrets are respectively defined as, $R(n):=\sum_{i=1}^{n}\big{(}\theta_{a_{i}^{*},i}-\theta_{a_{i},i}\big{)}$, $R^{(e)}(n):=\sum_{i=1}^{n}\big{(}\theta_{a_{i}^{(e)*},i}^{(e)}-\theta_{a_{i},i}^{(e)}\big{)}$ and $R^{(s)}(n):=\sum_{i=1}^{n}\big{(}\theta_{a_{i}^{(s)*},i}^{(s)}-\theta_{a_{i},i}^{(s)}\big{)}$, where $\theta_{a,i}:=\bm{x}_{i}^{T}[w\bm{\beta}_{a}^{*}+(1-w)\bm{\gamma}_{a}^{*}]$, $\theta_{a,i}^{(e)}:=\bm{x}_{i}^{T}\bm{\beta}_{a}^{*}$ and $\theta_{a,i}^{(s)}:=\bm{x}_{i}^{T}\bm{\gamma}_{a}^{*}$.
Figure 2: Box plots of cumulative regret based on three different criteria: utility, efficacy, and safety, at different stages of trials across $n_{sim}$ replications. Panels are arranged in columns based on three criteria, and two different rows correspond to the High-SNR and Low-SNR setups. The y-limits are set at different ranges across two different rows in order to compare cumulative regrets within each data-generating setup. A lower value of regret indicates better performance.
Results
Figure 2 presents three different types of cumulative regret incurred by all methods. The variability of regrets is similar within each data-generating setup, up to a different scaling. Lower regret is better as it indicates allocation of better treatments. Rand naturally has high cumulative regrets in all scenarios. TS optimizes arm allocation based solely on efficacy, resulting in the lowest cumulative efficacy regret but an even higher cumulative safety regret than Rand. Note that RiTS achieves the lowest cumulative safety regret and comparable utility regret as compared to TS. Although RiTS incurs higher efficacy regret than TS. Appendix A.3 also compares methods when functions $\{\mu^{(a)}(\cdot)\}_{a=1}^{K},\ \{\nu^{(a)}(\cdot)\}_{a=1}^{K}$ in (1) are mis-specified.
Figure 3: Left panels: Box plot for the frequency of arm allocations by different methods across $n_{sim}$ replications. Right panels: Panels in the first row present the proportion of replication where the stopping criteria are met across different stages of the trial. Panels in the second row depict the proportion replications where Arm 4 has the highest estimated effect size.
Left panels in Figure 3 depict frequency of arms allocated by different methods. TS allocates Arm 4 to more participants in the majority of trials, disregarding its lower safety profile. In contrast, RiTS allocates many participants to Arm 4 but also prioritizes Arm 3 for many due to Arm 3’s higher utility. Right panels of Figure 3 summarize performance based on two metrics introduced before. For both SC and WA4, a higher value indicates better performance. RiTS-AIPW meets expectation and performs the best in identifying Arm 4 as the best arm (i.e., WA4 metric) for both SNR setups. Both Rand-AIPW and TS-AIPW demonstrate comparable performance in terms of WA4. Despite its widespread use, Rand-T-test’s performance is slightly worse than previous methods. Rand-IPW, TS-IPW and RiTS-IPW consistently show poor performance throughout trials. Similar inconsistency carried over to the SC metric, where all IPW based AsympCS are observed to be not suitable for practical use in our setup. To save space, we exclude all IPW based methods from our discussions; however included in tables and figures for comparisons. Note that Rand-T-test leads to the highest number of abruptly stopped trials; it should be ignored due to inflated Type-1 error. During earlier stages, RiTS-AIPW shows comparable performance with Rand-AIPW to stop trials early with some confidence (i.e., SC metric). During later stages, Rand-AIPW leads to the largest number of abruptly stopped trials. RiTS-AIPW fares comparably with Rand-AIPW in High-SNR. For Low-SNR, RiTS-AIPW shows a slight drop in performance. TS-AIPW’s performance is comparable with RiTS-AIPW and Rand-AIPW in High-SNR. However, for Low-SNR, TS-AIPW leads to fewer abruptly stopped trials.
In order to gain further insights, we investigate performances based on finer metrics such as width, bias, root mean squared error (RMSE), and cumulative miscoverage of AsympCS. Web Appendices A.1 and A.2 include an elaborate discussion on this direction. We summarize our findings here. Rand-T-test yields the narrowest confidence intervals, however suffers from inflated cumulative mis-coverages. Rand-AIPW leads to the narrowest confidence sequences (CS) with valid time-uniform coverage. RiTS-AIPW yields slightly wider CS compared to Rand-AIPW. CSs from TS-AIPW are the widest, except for $\Delta(4)$, where widths of TS-AIPW are comparable to RiTS-AIPW. Narrower CSs lead to more abruptly stopped trials, which explains our previous discussions on the performance of these methods based on SC metric. Widths of CSs for IPW based AsympCS are much larger.
Setup Arm Method 50 60 70 80 90 100 110 120 150 200
High-SNR Arm2 Rand-T-Test 0(0.17) 0(0.15) 0(0.14) 0(0.13) 0(0.12) 0(0.12) 0(0.11) 0(0.11) 0(0.09) 0(0.08)
Rand-AIPW 0.01(0.12) 0(0.11) 0(0.1) 0(0.1) 0(0.09) 0(0.09) 0(0.08) 0(0.08) 0(0.07) 0(0.06)
TS-AIPW 0(0.16) 0(0.15) 0(0.15) 0(0.14) 0(0.13) 0(0.13) 0(0.12) 0(0.12) 0(0.11) 0(0.09)
RiTS-AIPW 0(0.14) 0(0.13) 0(0.12) 0(0.11) 0(0.11) 0(0.11) 0(0.1) 0(0.1) 0(0.09) 0(0.08)
Rand-IPW -0.33(0.37) 0(0.25) 0.01(0.34) 0(0.38) 0(0.39) 0(0.4) -0.01(0.4) -0.01(0.39) 0(0.38) -0.01(0.36)
TS-IPW 0(0.89) 0(0.88) 0.01(0.86) 0.01(0.84) -0.01(0.82) -0.02(0.8) -0.04(0.78) -0.03(0.77) -0.02(0.7) -0.03(0.62)
RiTS-IPW 0.01(0.72) 0(0.76) 0.01(0.74) 0(0.71) -0.02(0.69) -0.04(0.69) -0.03(0.68) -0.03(0.66) -0.03(0.62) -0.03(0.56)
Arm3 Rand-T-Test 0(0.09) 0(0.08) 0(0.08) 0(0.07) 0(0.07) 0(0.06) 0(0.06) 0(0.06) 0(0.05) 0(0.05)
Rand-AIPW 0.01(0.09) 0(0.09) 0(0.08) 0(0.08) 0(0.07) 0(0.07) 0(0.06) 0(0.06) 0(0.05) 0(0.05)
TS-AIPW 0(0.12) 0(0.11) 0(0.1) 0(0.1) 0(0.09) 0(0.09) 0(0.08) 0(0.08) 0(0.07) 0(0.06)
RiTS-AIPW 0(0.11) 0(0.1) 0(0.1) 0(0.09) 0(0.09) 0(0.08) 0(0.08) 0(0.07) 0(0.07) 0(0.06)
Rand-IPW -0.34(0.35) 0(0.22) 0.01(0.34) 0(0.38) 0.01(0.4) 0.02(0.41) 0.01(0.42) 0.01(0.41) 0.01(0.4) 0(0.37)
TS-IPW -0.01(0.87) -0.01(0.89) 0.01(0.85) -0.01(0.85) 0(0.83) -0.01(0.8) -0.03(0.78) -0.03(0.76) -0.01(0.73) -0.02(0.63)
RiTS-IPW -0.04(0.7) -0.06(0.69) -0.06(0.67) -0.04(0.66) -0.04(0.65) -0.05(0.65) -0.04(0.62) -0.05(0.61) -0.04(0.58) -0.03(0.51)
Arm4 Rand-T-Test -0.01(0.17) -0.01(0.15) 0(0.14) 0(0.13) 0(0.13) 0(0.12) 0(0.12) 0(0.11) 0(0.1) 0(0.09)
Rand-AIPW 0.01(0.12) 0(0.11) 0(0.1) 0(0.09) 0(0.09) 0(0.08) 0(0.08) 0(0.08) 0(0.07) 0(0.06)
TS-AIPW 0(0.14) 0(0.14) 0(0.13) 0(0.12) 0(0.11) 0(0.11) 0(0.11) 0(0.1) 0(0.09) 0(0.08)
RiTS-AIPW 0(0.14) 0(0.14) 0.01(0.13) 0(0.12) 0(0.11) 0(0.11) 0(0.11) 0(0.1) 0(0.09) 0(0.08)
Rand-IPW -0.35(0.39) -0.01(0.26) -0.01(0.37) -0.02(0.42) -0.03(0.44) -0.03(0.45) -0.04(0.45) -0.03(0.46) -0.04(0.44) -0.04(0.41)
TS-IPW 0.01(0.73) 0(0.73) 0(0.7) 0(0.68) -0.01(0.66) -0.02(0.63) -0.03(0.61) -0.04(0.58) -0.02(0.54) -0.03(0.49)
RiTS-IPW 0(0.71) -0.02(0.72) -0.02(0.7) 0(0.67) 0(0.67) -0.03(0.67) -0.03(0.63) -0.02(0.6) -0.03(0.56) -0.02(0.49)
Low-SNR Arm2 Rand-T-Test 0(0.09) 0(0.08) 0(0.08) 0(0.07) 0(0.07) 0(0.06) 0(0.06) 0(0.06) 0(0.05) 0(0.04)
Rand-AIPW 0.01(0.07) 0(0.07) 0(0.06) 0(0.06) 0(0.05) 0(0.05) 0(0.05) 0(0.05) 0(0.04) 0(0.04)
TS-AIPW 0(0.09) 0(0.09) 0(0.08) 0(0.08) 0(0.08) 0(0.08) 0(0.07) 0(0.07) 0(0.06) 0(0.05)
RiTS-AIPW 0(0.08) 0(0.07) 0(0.07) 0(0.06) 0(0.06) 0(0.06) 0(0.06) 0(0.06) 0(0.05) 0(0.04)
Rand-IPW -0.16(0.19) 0(0.13) 0(0.17) 0(0.19) 0(0.2) 0(0.2) -0.01(0.2) 0(0.2) 0(0.19) -0.01(0.18)
TS-IPW -0.01(0.44) 0(0.43) 0(0.43) -0.01(0.42) -0.02(0.4) -0.01(0.38) -0.01(0.38) -0.01(0.37) 0(0.34) 0(0.29)
RiTS-IPW 0.01(0.34) 0.01(0.33) 0.01(0.33) 0(0.31) 0(0.31) 0(0.3) 0(0.3) 0(0.29) 0(0.26) 0(0.24)
Arm3 Rand-T-Test 0(0.06) 0(0.05) 0(0.05) 0(0.05) 0(0.04) 0(0.04) 0(0.04) 0(0.04) 0(0.03) 0(0.03)
Rand-AIPW 0.01(0.06) 0(0.05) 0(0.05) 0(0.04) 0(0.04) 0(0.04) 0(0.04) 0(0.04) 0(0.03) 0(0.03)
TS-AIPW 0(0.07) 0(0.07) 0(0.06) 0(0.06) 0(0.06) 0(0.05) 0(0.05) 0(0.05) 0(0.05) 0(0.04)
RiTS-AIPW 0(0.06) 0(0.06) 0(0.06) 0(0.05) 0(0.05) 0(0.05) 0(0.05) 0(0.05) 0(0.04) 0(0.04)
Rand-IPW -0.17(0.18) 0(0.11) 0.01(0.17) 0(0.19) 0(0.2) 0.01(0.21) 0.01(0.21) 0(0.2) 0(0.2) 0(0.19)
TS-IPW -0.02(0.42) -0.02(0.41) 0(0.41) -0.01(0.41) -0.01(0.4) -0.01(0.39) -0.01(0.38) -0.02(0.37) -0.01(0.36) 0(0.32)
RiTS-IPW -0.01(0.34) -0.01(0.34) -0.01(0.33) -0.01(0.32) 0(0.31) -0.01(0.3) -0.01(0.3) 0(0.29) -0.01(0.26) -0.01(0.24)
Arm4 Rand-T-Test 0(0.09) 0(0.08) 0(0.08) 0(0.07) 0(0.07) 0(0.06) 0(0.06) 0(0.06) 0(0.05) 0(0.05)
Rand-AIPW 0.01(0.07) 0(0.06) 0(0.06) 0(0.05) 0(0.05) 0(0.05) 0(0.05) 0(0.05) 0(0.04) 0(0.04)
TS-AIPW 0(0.08) 0(0.08) 0(0.07) 0(0.07) 0(0.07) 0(0.06) 0(0.06) 0(0.06) 0(0.05) 0(0.05)
RiTS-AIPW 0(0.08) 0(0.08) 0(0.07) 0(0.07) 0(0.06) 0(0.06) 0(0.06) 0(0.06) 0(0.05) 0(0.05)
Rand-IPW -0.18(0.2) -0.01(0.13) -0.01(0.19) -0.01(0.21) -0.01(0.22) -0.02(0.22) -0.02(0.23) -0.02(0.23) -0.02(0.22) -0.02(0.2)
TS-IPW 0(0.37) 0(0.36) 0(0.35) 0(0.34) 0(0.32) -0.01(0.31) 0(0.3) -0.01(0.29) 0(0.27) -0.01(0.24)
RiTS-IPW 0(0.35) 0.01(0.34) 0.01(0.34) 0.01(0.33) 0.01(0.33) 0.01(0.32) 0.01(0.3) 0.01(0.29) 0.01(0.27) 0(0.25)
Table 1: Average bias (and RMSE) of estimated effect sizes based on different methods across 1000 replications. The first and last twelve rows correspond to High-SNR and Low-SNR setups. Different columns refer to the number of observations used to obtain estimates at different stages of trials. Our proposed method is RiTS-AIPW. Lower average bias and RMSE indicate better estimation.
Table 1 presents the average bias and RMSE of estimated effect sizes of all active doses. Bias for Low-SNR is approximately half the ones for High-SNR setup due to the change in true average efficacy. Performance of RiTS-AIPW is comparable to Rand-AIPW and TS-AIPW, although Rand-AIPW has slightly lower RMSE than both adaptive methods. Bias of Rand-T-test is slightly higher in early stages, however comparable to other methods in later stages. This explains the slight dip in Rand-T-test’s performance based on WA4 metric discussed before. Web Appendix A.2 further explores the validity of AsympCS empirically, for different choices of hyperparameters. Rand-AIPW successfully controls cumulative miscoverage in all scenarios. For adaptive methods, burn-in sample size $m$, number of initially randomized participants $n_{0}$, and clipping $\delta$ play a crucial role in controlling the cumulative miscoverage rate. We recommend fixing $\delta=0.1$ to encourage adaptive methods to explore sub-optimal arms sufficiently. A larger value of $m$ is also crucial to allow AsympCS to enter its asymptotic regime. Web Tables 1 and 3 use $m=80$. Larger values of $n_{0},m,\ \delta$ allow adaptive methods to allocate enough sample sizes across all arms, so that Assumption 1 is satisfied. RiTS-AIPW demonstrates nominal cumulative coverage when $\delta$ is set to 0.10.10.1. In other scenarios, TS-AIPW fails to provide nominal cumulative coverage when adaptivity is higher, e.g., in High-SNR. RiTS-AIPW also fails in some of those cases, but its performance is significantly better than TS-AIPW in controlling the cumulative miscoverage.
We now summarize our insights from the experiments and provide our recommendations to implement this framework in practice: (1) RiTS-AIPW is a much safer alternative compared to Rand-AIPW and TS-AIPW. (2) Severely mis-specified trial optimizers can potentially impact allocation of better and safer treatments, so domain knowledge from earlier phases has to be incorporated to mitigate mis-specifications. (3) Setting $\delta=0.05\ \text{or}\ 0.1$, $n_{0}=24\ \text{or}\ 40$, and $m=80$ seems sensible in scenarios similar to our settings. (4) A sufficiently large burn-in sample size $m$ is required for valid time-uniform coverage. During earlier stages, point estimates of RiTS-AIPW can be used safely. With enough burn-in sample size $m$, bounds in RiTS-AIPW can be used safety to abruptly stop trials with confidence.
4 Application to Phase IIb Trial for Treating Alopecia Areata
Alopecia areata is an autoimmune disorder characterized by sudden hair loss that can affect any hair-bearing area of the body. This condition often has significant psychological and emotional impacts, leading to decreased self-esteem and quality of life. Finding a treatment for alopecia areata is essential, as current treatment options, such as corticosteroids and immunosuppressants, are inadequate. A drug targeting the underlying autoimmune mechanism could potentially restore hair growth, improving patients’ well-being and quality of life. Thus, continued research efforts toward developing effective treatments are crucial to addressing the unmet medical needs of individuals with alopecia areata.
We use a real dataset from a recently conducted Phase IIb/III trial for a drug to treat alopecia areata. The study is a randomized trial involving five doses of varying levels with a placebo control. Phase IIb is continued to Phase III seamlessly without any final analysis at the end of Phase IIb. We use the percentage change in SALT score from baseline as the efficacy endpoint, which measures the percentage of scalp hair loss. For safety, we use the percentage change in lymphocyte counts in cells per microliter of blood from baseline. We consider the endpoints at Week 24, which is the optimal time for the drug to show its potential effects. Based on related research from the trial, we select baseline SALT score, baseline lymphocyte counts, age, BMI, albumin, and lymphocyte marker as potentially useful covariates to model the efficacy and safety endpoints. A significant challenge in applying bandit algorithms to historical datasets is the lack of knowledge about counterfactual endpoints for doses not allocated to each participant. Since it is a randomized trial, we use random forests (Breiman,, 2001) to estimate counterfactual endpoints based on a comprehensive set of demographic covariates and baseline lab measurements. The dataset includes 654 participants. We perform 1000 replications, where each replication involves sampling 654 participants with replacement, followed by applying one of Rand, TS, and RiTS. Whenever the allocated arm from the bandit does not match the allocation, we use the counterfactual efficacy and safety; otherwise, we use the observed endpoint values. Initially, arms are randomized to $n_{0}=60$ participants. The clipping parameter $\delta$ is set to 0.050.050.05. All other tuning parameters are the same as in Section 3. Figure 4 compares and summarizes the performance of all methods.
Figure 4: Left panel: Box plot for the frequency of arm allocations by different methods across 100010001000 replications. Right panel: Left sub-figure presents the proportion of replication where the stopping criteria are met across different stages of the trial, with the minimum clinically significant effect size set to 0.1. The right sub-figure depicts the proportion of replications where Arm 6 has the highest estimated effect size.
Arm 1 is set as the placebo, and Arms 2–6 represent increasing dose levels. Average efficacy of Arms 1–6 are 0.52-0.52- 0.52, 0.63-0.63- 0.63, 0.01-0.01- 0.01, 0.310.310.31, 0.220.220.22, and 0.440.440.44, whereas average safety of Arms 1–6 are 0.460.460.46, 0.130.130.13, 0.06-0.06- 0.06, 0.11-0.11- 0.11, 0.07-0.07- 0.07, and 0.30-0.30- 0.30, respectively. With $w=0.5$, Arm 4 has the highest utility, and we expect RiTS (with $w=0.5$) to allocate more participants to Arm 4. From the left panel of Figure 4 we observe the following: TS is more aggressive than RiTS, allocating Arms 1 and 2 to almost no participants after initial exploration. TS prioritizes Arm 6, focusing on higher efficacy while disregarding the safety profile. RiTS continues to allocate Arm 1 or Arm 2 to participants, even at later stages. Notably, RiTS allocates Arm 6 to fewer participants and instead slightly favors Arms 4 and 5 compared to Arm 6 due to the former’s better safety profiles. The right panel of Figure 4 highlights the benefit of RiTS over TS due to former’s balanced allocation. RiTS-AIPW recovers Arm 6 as the best arm more often in earlier stages. It also met the stopping criteria more often compared to TS-AIPW. Naturally, Rand-AIPW performs the best in both aspects due to an almost equal allocation of participants across all arms. Arm 4 was approved for commercial use in the study due to its favorable safety profile, while Arms 5 and 6 were considered potential candidates for approval. While efficacy is commonly prioritized for selecting the optimal dose, safety is also considered during the selection process. Construction of AsympCS can be naturally extended for average utility (for some $w$) instead of average efficacy. We focus on average efficacy for simplicity to illustrate our idea. Our empirical results align with the truth that Arm 6 was the most efficacious arm, while Arm 4 has the highest utility and therefore is favored most during adaptive randomization by RiTS. Good empirical performance in the real data provides further evidence that RiTS is a safer and more practical option for applying contextual bandit in clinical trials, a key step towards expediting drug development pipelines.
5 Discussion
This article proposes a novel method of response adaptive randomization (RAR) called Risk-inclusive Thompson sampling (RiTS) in dose-ranging studies. Our method offers a flexible approach to balance efficacy and safety while leveraging participant-level covariates during arm allocation. We also extend Asymptotic Confidence Sequences (AsympCS, Waudby-Smith et al.,, 2024) to perform time-uniform inference on effect sizes. Our approach is model-assisted: arm allocation relies on parametric assumptions; however, inference of effect sizes is robust to model mis-specifications. Both simulation experiments and application on a real-world dataset provide empirical evidence that our method is safer compared to traditional randomized trials.
There are several promising directions for future research. First, adaptive designs introduce additional operational challenges. For instance, we have explored a scenario of delayed response in our simulations. In the real data, we have access to endpoints across multiple follow-ups of each participant. It would be interesting to utilize observed endpoints in early follow-ups for adaptive allocation of arms. Moreover, RAR methods often unblind clinical trial staff to results of the trial due to strings of treatment assignments to one arm (Proschan and Evans,, 2020). Further research is necessary to address these issues. Second, extending contextual bandits and AsympCS to accommodate different endpoint types, such as binary, count, or time-to-event, is of independent interest; see Dumitrascu et al., (2018) and references therein. Third, regarding inference, asymptotic approximation requires an adequate burn-in sample size, which may be a key limitation in early-phase trials. Developing alternative anytime valid inferential strategies is of independent interest.
Finally, a fundamental trade-off exists between efficiently estimating effect sizes and ensuring participant safety in clinical trials. ER optimizes estimation but compromises trial safety, whereas nonrandomized (deterministic) procedures exploit the best treatment, while safer, but introduces bias. RAR methods, such as RiTS, aim to balance these extremes. Further research is needed to investigate how to balance these two extremes to conduct safer, shorter, and more efficient trials.
Acknowledgments
Part of the research was conducted during Kanrar’s 2024 Summer Internship at Pfizer in collaboration with Ghodsi and Gamalo. The authors thank Professor Philip Dixon for helpful comments on a previous version of this manuscript. Kanrar also thanks Professor Somak Dutta for providing access to his personal High-performance Computing cluster. The authors acknowledge the use of generative AI tools to assist in refining the language of the manuscript.
Supporting Information
Reproducible R code for all experiments is made publicly available at https://github.com/rohitkanrar/RiTS.
Data Availability
Data sharing is not applicable to this article, as no new datasets are generated. The method developed will enable the design and analysis of future early-phase clinical trials.